# Hybrid Beamforming Design for RSMA-assisted mmWave Integrated Sensing and Communications

### v1

@@ Abstract
Integrated sensing and communications (ISAC) has been considered one of the new paradigms for sixth-generation (6G) wireless networks. In the millimeter-wave (mmWave) ISAC system, hybrid beamforming (HBF) is considered an emerging technology to exploit the limited number of radio frequency (RF) chains in order to reduce the system hardware cost and power consumption. However, the HBF structure reduces the spatial degrees of freedom for the ISAC system, which further leads to increased interference between multiple users and between users and radar sensing. To solve the above problem, rate split multiple access (RSMA), which is a flexible and robust interference management strategy, is considered. We investigate the joint common rate allocation and HBF design problem for the HBF-based RSMA-assisted mmWave ISAC scheme. We propose the penalty dual decomposition (PDD) method coupled with the weighted mean squared error (WMMSE) minimization method to solve this high-dimensional non-convex problem, which converges to the Karush-Kuhn-Tucker (KKT) point of the original problem. Then, we extend the proposed algorithm to the HBF design based on finite-resolution phase shifters (PSs) to further improve the energy efficiency of the system. Simulation results demonstrate the effectiveness of the proposed algorithm and show that the RSMA-ISAC scheme outperforms other benchmark schemes.
@@ I Introduction
The sixth generation (6G) of wireless networks is regarded as a major enabler for emerging services such as autonomous driving, smart manufacturing, remote health care, and digital twins [1]. The aforementioned applications require networks with both high-throughput communication and high-accuracy sensing capabilities. Meanwhile, the exponential growth of wireless services and the massive number of connections to wireless devices are straining spectrum resources, which will create an urgent need for additional spectrum, of which the radar frequency bands are one of the best candidates for coexistence with communication systems [2]. Based on the above service and spectrum coexistence requirements, integrated sensing and communications (ISAC) technology has arisen and attracted extensive resesarch attention in industry and academia [3]. Recently, the international telecommunication union (ITU) has released the 6G overall objective recommendation, in which ISAC, as one of the six key scenarios, aims to enable ubiquitous wireless sensing capabilities for 6G network infrastructures [4].
Distinguished from communication and radar coexistence designs [5],[6], ISAC is designed to simultaneously realize communication and sensing functions by sharing wireless resources, hardware platforms, and joint signal processing frameworks, yielding integration gains and collaboration gains to the system. Specifically, ISAC signal design methods can be classified into three categories: sensing-centric design (SCD), communication-centric design (CCD), and joint design (JD). Early SCD scheme modulates the communication symbols into the radar chirp signals [7]. The CCD scheme is mainly based on orthogonal frequency division multiplexing (OFDM) waveforms to design the ISAC signal [8], [9]. However, the JD scheme does not rely on the existing waveforms and designs the ISAC waveforms from scratch in order to improve the freedom and flexibility of the system. The authors of [10] pioneered a JD approach based on optimization techniques to solve the problem of designing the constant envelope waveforms for the ISAC system. In addition, since multiple-input multiple-output (MIMO) technology can further improve the diversity gain and degrees of freedom (DoFs) of the system, the integration design of MIMO communication and MIMO radar in the spatial domain is an inevitable trend. The authors of [11] proposed a scheme that utilizes the sidelobe level of the MIMO radar beampattern to represent the communication symbols, while the main beam is only used for target sensing. Further, the beamforming design of the MIMO-ISAC system can be formulated as nonlinear optimization problems considering communication performance metrics (e.g., throughput [12], [13], symbol error rate (SER) [14], singal-to-interference-plus-noise ratio (SINR) [15], etc.) and sensing performance metrics (e.g., beampattern matching mean-square error [16],[17], signal-to-clutter-plus-noise ratio (SCNR) [18], Cramér-Rao bound (CRB) [19], [20], etc.).
All of the above ISAC research is based on the sub-6 GHz band, but in order to cope with the explosive growth of wireless devices and services, the millimeter-wave (mmWave) spectrum band of 30-300 GHz can further dramatically improve communication capacity and target sensing accuracy, and thus, become a new research spectrum band for ISAC. The authors of [21] achieved gigabit communication rates and centimeter-level sensing accuracy for short distances based on the IEEE 802.11ad protocol at 60 GHz carriers. However, the short wavelength of mmWave leads to severe road loss and rain fading [22]. Therefore, transmitters usually need to employ massive MIMO (mMIMO) technology to form high beam gains to overcome these drawbacks. Note that the fully digital large-scale antenna array structure, i.e., each antenna needs to be equipped with a radio frequency (RF) chain, will lead to high hardware costs and power consumption. To address the above issues, the mmWave communication base station (BS) has widely adopted hybrid beamforming (HBF) architectures [23],[24],[25], which require only a few RF chains to be connected to the large-scale antenna arrays via phase shifters (PSs), thus reducing the hardware cost and improving the energy efficiency of the system. In addition, the above hybrid architecture was also demonstrated in phased-MIMO radar [26], which integrates the advantages of the diversity gain of orthogonal waveforms of MIMO radar and the high beam gain of phased-array radar to improve the SCNR of the received signals and the target detection probability. Based on the similarities between communication hybrid architectures and phased-MIMO radars, the research of HBF has been in mmWave ISAC systems. The authors of [27] have pioneered the HBF design for ISAC systems that minimizes the weighted sum of the radar beampattern error and the communication precoding error. The complete uplink and downlink communication protocols and target sensing steps in HBF-based mmWave ISAC systems were given in [28]. The authors of [29],[30] presented the HBF design scheme based on the alternating direction method of multipliers (ADMM) algorithm.
@@
In the MIMO-ISAC systems mentioned above, spatial division multiple access (SDMA) was used to manage the interference between the user and the radar. However, rate-splitting multiple access (RSMA) is a non-orthogonal transmission strategy based on rate-splitting (RS) precoding at the transmitter and successive interference cancellation (SIC) decoding at the receiver, which provides strong anti-interference capability for multi-antenna wireless networks [31]. Specifically, RSMA divides user messages into common and private parts, encodes the common part jointly into common data streams decoded by SIC for multiple users, and encodes the private part independently into private data streams decoded by corresponding users. RSMA reconciles two extreme interference management strategies, SDMA, which treats interference entirely as noise, and non-orthogonal multiple access (NOMA), which decodes interference completely, to achieve higher spectral and energy efficiencies [32],[33]. Further, the authors of [34],[35] proposed a scheme for the RSMA-assisted ISAC system and demonstrated the advantages of RSMA in managing interference between communication users and interference between dual functions. The authors of [36] investigated RSMA-assisted multi-BS cooperative ISAC systems for advanced interference management and high-accuracy localization services.
@@
Motivated by the advantages of mmWave hybrid architecture‘s low power consumption and robust RSMA interference management, we introduce RSMA in mmWave ISAC systems to address the joint common rate allocation and HBF design issues in order to fill the research gap in this area. We further improve the performance trade-off region of the ISAC system and reveal the mechanism of the RSMA strategy common data streams at the dual functions of communication and sensing. The main contributions of this paper are summarized as follows:
  • 1)
    We first propose mmWave ISAC system with HBF structure assisted by RSMA. We formulate the weighted sum rate (WSR) maximization problem by joint common rate allocation and HBF design, subject to the constraints of target sensing SCNR threshold, PSs unit-modulus, the total power budget, and the necessity for accurate correct decoding of common messages.
  • 2)
    The penalty dual decomposition (PDD) method coupled with the weighted mean squared error minimization (WMMSE) method is proposed to solve the above high-dimensional non-convex problem. First, the WMMSE method is used to transform the original problem into an equivalent form that is easy to solve. Next, the PDD algorithm of [37],[38] is used to solve the corresponding problem by means of an inner and outer double loop, in which the variables are updated in the form of a closed-form solution in the inner loop. Rigorous proof of the algorithm’s converges to the Karush-Kuhn-Tucker (KKT) point of the original optimization problem is provided.
  • 3)
    We extend the proposed algorithm to the HBF design based on finite-resolution PSs, which further reduces the hardware cost and power consumption and improves the energy efficiency of the system.
  • 4)
    Our simulation results show that the proposed WMMSE-PDD algorithm converges quickly. For both high and low spatial correlation channels, the RSMA-ISAC scheme outperforms the SDMA-ISAC and NOMA-ISAC schemes in terms of WSR and anti-clutter interference capability. We reveal that the common data stream has a dual function in mitigating multi-user interference and trade-offs between communication and sensing performance.
@@
The remainder of this paper is organized as follows. In Section II, we introduce the HBF-based RSMA-assisted mmWave ISAC system model and problem formulation. The specific algorithm for solving the optimization problem is presented in Section III. Section IV presents the extension of the algorithm to ISAC systems with finite resolution PSs. The simulation results are presented in Section V. Finally, our conclusions are provided in Section VI.
Notations: Upper-case and lower-case boldface letters denote matrices and vectors, respectively. $(\cdot)^{*}$, $(\cdot)^{T}$, $(\cdot)^{H}$, $(\cdot)^{-1}$, and $(\cdot)^{\dagger}$ denote complex conjugate, transpose, Hermitian transpose, matrix inversion, and Moore-Penrose matrix inversion, respectively. $\operatorname{tr}(\cdot)$, $\left\|\cdot\right\|_{p}$, and ${}\Re\{\cdot\}$ denote the trace of the matrix, $l_{p}$ norm operation, and the real part of a complex variable, respectively. $\mathbb{E}\{\cdot\}$ denotes the expectation. $\mathbb{C}^{m\times n}$ denotes the space of $m\times n$ complex matrices. $\left|\boldsymbol{\rm A}(m,n)\right|$ denotes the modulus of the element in the $m$-th row and $n$-th column of matrix A. ${\rm vec\{\cdot\}}$ denotes the matrix vectorization. ${\rm diag}\{\boldsymbol{\rm c}\}$ denotes the diagonal matrix, whose diagonal elements are formed by the vector $\boldsymbol{\rm c}$. $\otimes$ denotes the Kronecker products between two matrices.
Figure 1: RSMA-assisted mmWave ISAC system.
@@ II System Model and Problem Formulation
In this section, we first introduce the RSMA-assisted mmWave ISAC system model, the performance metrics of the communication, and sensing functions, and then formulate the HBF problem.
II-A Transmit Model
@@
Consider an RSMA-assisted mmWave ISAC system as shown in Fig. 1, where one BS with $N_{t}$ transmit antennas and $N_{r}$ receive antennas simultaneously serve $K$ single-antenna downlink user communications while engaging in radar target detection. Let $\mathcal{K}=\{1,2,\cdots,K\}$ represent the set of all user indexes. Specifically, the messages of $K$ users are first split and encoded to form $K+1$ data streams via the RSMA technique. Further, to reduce the hardware cost and complexity, we use an HBF structure to transmit the data stream, which is first digitally processed at the baseband using a digital beamformer, and then the digitally processed signal is upconverted to the carrier frequency through the RF chains, followed by PSs to implement an analog beamformer. Without loss of generality, it is assumed that the multiple antennas are uniform linear arrays (ULA) and the number of RF chains is ${N}_{\rm RF}$ with the constraint $K+1\leq{N}_{\rm RF}\leq N_{t}$. Since a monostatic sensing model is considered, the transmitter operates in time-division duplex (TDD) mode to avoid interference between the transmit and sensing echo signals.
@@
According to [31], when we consider the RSMA transmission scheme, the message $W_{k}$ of the $k$-th user is first split into two parts, one of which is the common part $W_{c,k}$, and the other is the private part $W_{p,k}$, $\forall k\in\mathcal{K}$. The common data stream $s_{c}$ is obtained by encoding the common message $W_{c}$ formed by combining all user common parts $\{W_{c,1},\cdots,W_{c,K}\}$, while private data streams $\{s_{1},\cdots,s_{K}\}$ are obtained by encoding the private parts $\{W_{p,1},\cdots,W_{p,K}\}$. Define the vector $\boldsymbol{\rm s}=[s_{c},s_{1},\cdots,s_{K}]^{T}\in\mathbb{C}^{(K+1)\times 1}$ as data streams with unit power and satisfy $\mathbb{E}\{\boldsymbol{\rm s}\boldsymbol{\rm s}^{H}\}=\boldsymbol{\rm I}_{K+1}$. Next, the data stream $\boldsymbol{\rm s}$ undergoes hybrid beamforming to obtain a dual-function integrated signal that simultaneously achieves multi-user communication and radar target detection. Accordingly, the transmit ISAC signal, denoted by $\boldsymbol{\rm x}\in\mathbb{C}^{{N_{t}}\times 1}$, can be expressed as
$\boldsymbol{\rm x}$ $=\boldsymbol{\rm F}_{\rm RF}\boldsymbol{\rm F}_{\rm D}\boldsymbol{\rm s}$ (1)
$=\boldsymbol{\rm F}_{\rm RF}\boldsymbol{\rm f}_{c}s_{c}+\sum_{i\in\mathcal{K}}\boldsymbol{\rm F}_{\rm RF}\boldsymbol{\rm f}_{{\rm D},i}s_{i},$
@@
where $\boldsymbol{\rm F}_{\rm D}=[\boldsymbol{\rm f}_{c},\boldsymbol{\rm f}_{\rm D,1},\cdots,\boldsymbol{\rm f}_{{\rm D},K}]\in\mathbb{C}^{N_{\rm RF}\times(K+1)}$ is the digital beamforming matrix, in which $\boldsymbol{\rm f}_{c}$ is the common data stream precoder and $\boldsymbol{\rm f}_{{\rm D},k},k\in\mathcal{K}$ is the precoder for the $k$-th user’s private data stream. $\boldsymbol{\rm F}_{\rm RF}\in\mathbb{C}^{N_{t}\times N_{\rm RF}}$ is the analog beamforming matrix that is composed of $N_{t}\times N_{\rm RF}$ phase shifters with a fully connected structure [25]. As the phase shifter can only adjust the phase of the signal, all entries in the $\boldsymbol{\rm F}_{\rm RF}$ should satisfy the unit modulus constraints, e.g., $\left|\boldsymbol{\rm F}_{\rm RF}(i,j)\right|=1$.
II-B Communication Model And Metrics
@@
Under the above transmit signal model, the receive signal, denoted by $y_{k}$, at $k$-th single antenna user can be expressed as
$y_{k}$ $=\boldsymbol{\rm h}_{k}^{H}\boldsymbol{\rm x}+n_{k}$ (2)
$=\boldsymbol{\rm h}_{k}^{H}\boldsymbol{\rm F}_{\rm RF}\boldsymbol{\rm f}_{c}s_{c}+\sum_{i\in\mathcal{K}}\boldsymbol{\rm h}_{k}^{H}\boldsymbol{\rm F}_{\rm RF}\boldsymbol{\rm f}_{{\rm D},i}s_{i}+n_{k},$
@@
where ${\boldsymbol{\rm h}}_{k}\in\mathbb{C}^{N_{t}\times 1}$ is the channel vector from the transmitter to the $k$-th user and is assumed to be known by the BS. $n_{k}\sim\mathcal{CN}\left(0,\sigma_{{k}}^{2}\right)$,$k\in\mathcal{K}$, is the additive white Gaussian noise (AWGN) with zero mean and variance $\sigma_{{k}}^{2}$. Due to the sparse scattering environment of mmWave communications, the channel is directional and spatially sparse in the angular domain. Without loss of generality, we adopt the widely used Saleh-Valenzuela channel for mmWave communications [22], so $\boldsymbol{\rm h}_{k}$ can be written as
$\boldsymbol{\rm h}_{k}=\sqrt{\frac{N_{t}}{L_{k}}}\sum_{\ell=1}^{L_{k}}\beta_{\ell,k}\boldsymbol{\rm a}\left(\theta_{\ell,k}\right),$ (3)
where $L_{k}$ denotes the number of paths for the $k$-th user. $\beta_{1,k}$ and $\beta_{\ell,k},(2\leq\ell\leq L_{k})$ denote the complex gains of the $k$-th user line-of-sight (LoS) component and non-line-of-sight (NLoS) components, respectively. $\theta_{\ell,k}\in[-\frac{\pi}{2},\frac{\pi}{2}]$ is the angle of departure (AOD) of the $l$-th path, and the array steering vector $\mathbf{a}(\theta)\in\mathbb{C}^{N_{t}\times 1}$ can be given by
$\mathbf{a}(\theta)=\frac{1}{\sqrt{N_{t}}}\big{[}1,e^{j\frac{2\pi}{\lambda}d\sin(\theta)},\cdots,e^{j\left(N_{t}-1\right)\frac{2\pi}{\lambda}d\sin(\theta)}\big{]}^{T},$ (4)
where $\lambda$ and $d$ are the carrier wavelength and the antenna spacing, respectively. In this paper, we use a half-wavelength antenna array, i.e., $d=\lambda/2$.
According to the decoding criterion of the RSMA scheme [31], each user first decodes the common data stream $s_{c}$ into $\hat{W}_{c}$ by treating the interference from the private data streams $\{s_{1},\cdots,s_{K}\}$ as noise. Thus, the SINR, denoted by $\gamma_{c,k}$, of decoding $s_{c}$ at the $k$-th user is given by
${\gamma}_{c,k}=\frac{\left|\mathbf{h}_{k}^{H}\mathbf{F}_{\rm RF}\mathbf{f}_{c}\right|^{2}}{\sum\limits_{i\in\mathcal{K}}\left|\mathbf{h}_{k}^{H}\mathbf{F}_{\rm RF}\mathbf{f}_{{\rm D},i}\right|^{2}+\sigma_{{k}}^{2}}.$ (5)
@@
Next, the common data stream reconstructed by $\hat{W}_{c}$ is subtracted from the received signal $y_{k}$ using SIC. Then, user $k$ decodes the private data stream $s_{k}$ into $\hat{W}_{p,k}$ by treating the interference from private data streams $\{s_{i}\},i\neq k$ as noise. The SINR, denoted by $\gamma_{k}$, of decoding $s_{k}$ at the $k$-th user is then given by
@@
${\gamma}_{k}=\frac{\left|\mathbf{h}_{k}^{H}\mathbf{F}_{\rm RF}\mathbf{f}_{{\rm D},k}\right|^{2}}{\sum\limits_{i\in\mathcal{K},i\neq k}\left|\mathbf{h}_{k}^{H}\mathbf{F}_{\rm RF}\mathbf{f}_{{\rm D},i}\right|^{2}+\sigma_{{k}}^{2}}.$ (6)
@@
Finally, user $k$ splits $\hat{W}_{c,k}$ from the common message $\hat{W}_{c}$ and combines it with $\hat{W}_{p,k}$ to reconstruct the private message $\hat{W}_{k}$. Based on Eqs. (5) and (6), the achievable rates of decoding $s_{c}$ and $s_{k}$ for user $k$ are $R_{c,k}=\log_{2}\left(1+\mathrm{\gamma}_{c,k}\right)$ and $R_{k}=\log_{2}\left(1+\mathrm{\gamma}_{k}\right)$, respectively. To ensure that all users can successfully decode the common message, the rate of the common data stream $s_{c}$ can not exceed $R_{c}={\rm min}\{R_{c,1},\cdots,R_{c,K}\}$. According to the structure of RS, the rate of the common message is divided into $K$ portions, satisfying $\sum_{k\in\mathcal{K}}C_{k}=R_{c}$, where $C_{k}$ is the rate corresponding to the common part $W_{c,k}$ of the $k$-th user message. Thus, the total achievable rate for user $k$ consists of the sum of the rates of the common part $W_{c,k}$ and the private part $W_{p,k}$. In the RSMA-assisted ISAC system, for the communication function, we select the weighted sum rate (WSR), denoted by $R_{\rm WSR}$, as the performance metric of the system, which can be expressed as
$R_{\rm WSR}=\sum_{k\in\mathcal{K}}\alpha_{k}(C_{k}+R_{k}),$ (7)
where $\alpha_{k}$ is the weight factor representing the $k$-th user priority.
II-C Radar Sensing Model And Metrics

### v2

# {textblock*}

@@ Abstract
The harsh environment and scarce resources post-disaster drive the equipment to be miniaturized and portable. Based on this, integrated sensing and communication (ISAC) systems play a significant role in providing emergency wireless networks. In order to reduce the hardware cost, a hybrid beamforming (HBF) assisted millimeter-wave (mmWave) ISAC system, which exploits the limited number of radio frequency (RF) chains, is considered in this paper. However, the HBF structure reduces the spatial degrees of freedom, thus leading to increased interference among communication users and radar sensing. To solve this problem, a rate-splitting multiple access (RSMA) strategy is adopted to enhance the emergency mmWave-ISAC system. We formulate the weighted sum rate (WSR) maximization objective by jointly designing common rate allocation and HBF. Then, we propose the penalty dual decomposition (PDD) coupled with the weighted mean squared error (WMMSE) method to solve this high-dimensional non-convex problem. Numerical results demonstrate the effectiveness of the proposed algorithm and show that the RSMA-ISAC scheme outperforms other benchmark schemes.
@@ I Introduction
The rapid advancement of urbanization coupled with information construction, has inadvertently heightened the susceptibility of urban lifelines to disruptions in the wake of catastrophic events. After natural disasters, the fundamental infrastructure is often paralyzed[1]. Various items of emergency rescue equipment are taken with rescuers to the front line to meet the timing of disaster information acquisition and transmission. The harsh environment and scarce resources post-disaster drive the equipment to be integrated and portable[2]. To further increase the probability of survivals during the critical golden-rescue period[3], the aforementioned equipment must possess both high-throughput communication and high-accuracy sensing capabilities. However, the exponential growth of emergency services and the massive number of connections to rescue devices are straining spectrum resources, creating an urgent need for additional spectrum. Based on this, millimeter-wave (mmWave) based integrated sensing and communications (ISAC)[4], has been considered one of the new paradigms for emergency wireless networks[5].
@@
To compensate for the shortages of severe path-loss and rain attenuation in mmWave communications, massive multiple-input multiple-output (MIMO) is employed to generate high-gain directional beams[6][7]. To decrease the hardware costs and power consumption of fully digital large-scale antenna array structure, the hybrid beamforming (HBF) for ISAC systems was designed to connect the large-scale antenna arrays via phase shifters (PSs) with fewer radio frequency (RF) chains[8]. However, the adoption of HBF introduces challenges related to managing interference among communication users and radar sensing. Rate-splitting multiple access (RSMA) is a non-orthogonal transmission strategy based on rate-splitting (RS) precoding at the transmitter and successive interference cancellation (SIC) decoding at the receiver, which provides strong anti-interference capability for multi-antenna wireless networks [9]. RSMA reconciles two extreme interference management strategies, spatial division multiple access (SDMA), which treats interference entirely as noise, and non-orthogonal multiple access (NOMA), which decodes interference completely, to achieve higher spectrum and energy efficiencies. Further, the authors of [10] demonstrated the advantages of RSMA in managing interference between communication users and interference between dual functions of ISAC systems.
@@
Motivated by the advantages of mmWave hybrid architecture’s low power consumption and robust RSMA interference management, we introduce RSMA in mmWave ISAC systems to address the joint common rate allocation and HBF design issues to fill the research gap in this area. We further improve the performance trade-off region of the ISAC system and reveal the superiority of the RSMA common data streams at the dual functions of communication and sensing.
@@
The remainder of this paper is organized as follows. Section II gives the system model and problem formulation of RSMA-assisted emergency mmWave ISAC with HBF structure. The specific algorithm for solving the optimization problem is presented in Section III. The numerical results are presented in Section IV. Finally, our conclusions are provided in Section V.
@@ II System Model and Problem Formulation
In this section, we consider an RSMA-assisted emergency mmWave ISAC system as shown in Fig. 1, where one base station (BS) with $N_{t}$ transmit antennas and $N_{r}$ receive antennas simultaneously serve $K$ single-antenna downlink user communications while engaging in radar target detection. Let $\mathcal{K}=\{1,2,\cdots,K\}$ represent the set of all user indexes. Specifically, the messages of $K$ users are first split and encoded to form $K+1$ data streams via the RSMA technique. Without loss of generality, the number of RF chains is ${N}_{\rm RF}$ with the constraint $K+1\leq{N}_{\rm RF}\leq N_{t}$.
Figure 1: RSMA-assisted emergency mmWave ISAC system.
@@
Since the RSMA transmission scheme is adopted, the message $W_{k}$ of the $k$-th user is first split into two parts: one of which is the common part $W_{c,k}$, and the other is the private part $W_{p,k}$, $\forall k\in\mathcal{K}$. The common data stream $s_{c}$ is obtained by encoding the common message $W_{c}$ formed by combining all user common parts $\{W_{c,1},\cdots,W_{c,K}\}$, while private data streams $\{s_{1},\cdots,s_{K}\}$ are obtained by encoding the private parts $\{W_{p,1},\cdots,W_{p,K}\}$. Define the vector $\boldsymbol{\rm s}=[s_{c},s_{1},\cdots,s_{K}]^{T}\in\mathbb{C}^{(K+1)\times 1}$ as data streams with unit power and satisfy $\mathbb{E}\{\boldsymbol{\rm s}\boldsymbol{\rm s}^{H}\}=\boldsymbol{\rm I}_{K+1}$. Next, the data stream $\boldsymbol{\rm s}$ undergoes hybrid beamforming to obtain a dual-function integrated signal that simultaneously achieves multi-user communication and radar target detection. Accordingly, the transmit ISAC signal, denoted by $\boldsymbol{\rm x}\in\mathbb{C}^{{N_{t}}\times 1}$, can be expressed as
$\boldsymbol{\rm x}\!=\!\boldsymbol{\rm F}_{\rm RF}\boldsymbol{\rm F}_{\rm D}\boldsymbol{\rm s}\!=\!\boldsymbol{\rm F}_{\rm RF}\boldsymbol{\rm f}_{c}s_{c}\!+\!\sum_{i\in\mathcal{K}}\boldsymbol{\rm F}_{\rm RF}\boldsymbol{\rm f}_{{\rm D},i}s_{i},$ (1)
@@
where $\boldsymbol{\rm F}_{\rm D}=[\boldsymbol{\rm f}_{c},\boldsymbol{\rm f}_{\rm D,1},\cdots,\boldsymbol{\rm f}_{{\rm D},K}]\in\mathbb{C}^{N_{\rm RF}\times(K+1)}$ is the digital beamforming matrix, in which $\boldsymbol{\rm f}_{c}$ is the common data stream precoding vector and $\boldsymbol{\rm f}_{{\rm D},k},k\in\mathcal{K}$ is the precoding vector for the $k$-th user’s private data stream. $\boldsymbol{\rm F}_{\rm RF}\in\mathbb{C}^{N_{t}\times N_{\rm RF}}$ is the analog beamforming matrix that is composed of $N_{t}\!\times\!N_{\rm RF}$ phase shifters with a fully connected structure [11]. As the phase shifter can only adjust the phase of the signal, all entries in the $\boldsymbol{\rm F}_{\rm RF}$ should satisfy the unit modulus constraints, e.g., $\left|\boldsymbol{\rm F}_{\rm RF}(i,j)\right|=1$.
II-A Communication Model And Metrics
@@
Under the above transmit signal model, the received signal, denoted by $y_{k}$, at $k$-th single antenna user can be expressed as
$y_{k}=\boldsymbol{\rm h}_{k}^{H}\boldsymbol{\rm F}_{\rm RF}\boldsymbol{\rm f}_{c}s_{c}\!+\!\sum_{i\in\mathcal{K}}\boldsymbol{\rm h}_{k}^{H}\boldsymbol{\rm F}_{\rm RF}\boldsymbol{\rm f}_{{\rm D},i}s_{i}+n_{k},$ (2)
@@
where ${\boldsymbol{\rm h}}_{k}\in\mathbb{C}^{N_{t}\times 1}$ is the channel vector from the transmitter to the $k$-th user and is assumed to be known by the BS. $n_{k}\sim\mathcal{CN}\left(0,\sigma_{{k}}^{2}\right)$$k\in\mathcal{K}$, is the additive white Gaussian noise (AWGN) with zero mean and variance $\sigma_{{k}}^{2}$. Without loss of generality, we adopt the widely used Saleh-Valenzuela channel for mmWave communications [12], so $\boldsymbol{\rm h}_{k}$ can be written as $\boldsymbol{\rm h}_{k}=\sqrt{\frac{N_{t}}{L_{k}}}\sum_{\ell=1}^{L_{k}}\beta_{\ell,k}\boldsymbol{\rm a}\left(\theta_{\ell,k}\right)$, where $L_{k}$ is the number of channel paths for the $k$-th user. $\beta_{1,k}$ and $\beta_{\ell,k},(2\leq\!\ell\!\leq L_{k})$ denote the complex gains of the $k$-th user line-of-sight (LoS) component and non-line-of-sight (NLoS) components, respectively. $\theta_{\ell,k}\in[-\frac{\pi}{2},\frac{\pi}{2}]$ is the angle of departure (AOD) of the $l$-th path, and the array steering vector $\mathbf{a}(\theta)\in\mathbb{C}^{N_{t}\times 1}$ can be given by $\mathbf{a}(\theta)\!=\!\frac{1}{\sqrt{N_{t}}}\big{[}1,e^{j\frac{2\pi}{\lambda}d\sin(\theta)},\cdots,e^{j\left(N_{t}-1\right)\frac{2\pi}{\lambda}d\sin(\theta)}\big{]}^{T}$, where $\lambda$ and $d$ are the carrier wavelength and the antenna spacing, respectively. In this paper, we use a half-wavelength uniform linear array (ULA), i.e., $d=\lambda/2$.
@@
According to [9], the signal-to-interference-plus-noise (SINR) of decoding common data stream $s_{c}$ and private data stream $s_{k}$ at the $k$-th user, denoted by $\gamma_{c,k}$ and $\gamma_{k}$, respectively, can be given as follows:
@@
$\begin{cases}&{\gamma}_{c,k}=\frac{\left|\mathbf{h}_{k}^{H}\mathbf{F}_{\rm RF}\mathbf{f}_{c}\right|^{2}}{\sum\limits_{i\in\mathcal{K}}\left|\mathbf{h}_{k}^{H}\mathbf{F}_{\rm RF}\mathbf{f}_{{\rm D},i}\right|^{2}+\sigma_{{k}}^{2}}.\\ &{\gamma}_{k}=\frac{\left|\mathbf{h}_{k}^{H}\mathbf{F}_{\rm RF}\mathbf{f}_{{\rm D},k}\right|^{2}}{\sum\limits_{i\in\mathcal{K},i\neq k}\left|\mathbf{h}_{k}^{H}\mathbf{F}_{\rm RF}\mathbf{f}_{{\rm D},i}\right|^{2}+\sigma_{{k}}^{2}}.\end{cases}$ (3)
@@
Based on Eq.(3), the achievable rates of decoding $s_{c}$ and $s_{k}$ for user $k$ are $R_{c,k}=\log_{2}\left(1+\mathrm{\gamma}_{c,k}\right)$ and $R_{k}=\log_{2}\left(1+\mathrm{\gamma}_{k}\right)$, respectively. To ensure that all users can successfully decode the common message, the rate of the common data stream $s_{c}$ meets $R_{c}={\rm min}\{R_{c,1},\cdots,R_{c,K}\}$ and $\sum_{k\in\mathcal{K}}C_{k}=R_{c}$, where $C_{k}$ is the rate of the common stream $W_{c,k}$ intended for user- $k$. In the RSMA-assisted ISAC system, for the communication function, we select the weighted sum rate (WSR), denoted by $R_{\rm WSR}$, as the performance metric of the system, which can be expressed as $R_{\rm WSR}=\sum_{k\in\mathcal{K}}\alpha_{k}(C_{k}+R_{k})$, where $\alpha_{k}$ denotes the priority factor of the $k$-th user.
II-B Radar Sensing Model And Metrics